# Normalized Conditional Mutual Information Surrogate Loss for Deep Neural Classifiers

### v1

@@ Abstract
In this paper, we propose a novel information-theoretic surrogate loss—normalized conditional mutual information (NCMI)—as a drop-in alternative to the de facto cross-entropy (CE) for training deep neural network (DNN)-based classifiers. We first observe that the model’s NCMI is inversely proportional to its accuracy. Building on this insight, we introduce an alternating algorithm to efficiently minimize the NCMI. Across image recognition and whole-slide imaging (WSI) subtyping benchmarks, NCMI-trained models surpass state-of-the-art losses by substantial margins at a computational cost comparable to that of CE. Notably, on ImageNet, NCMI yields a 2.77% top-1 accuracy improvement with ResNet-50 comparing to the CE; on CAMELYON-17, replacing CE with NCMI improves the macro-F1 by 8.6% over the strongest baseline. Gains are consistent across various architectures and batch sizes, suggesting that NCMI is a practical and competitive alternative to CE. All code and data are publicly available at https://github.com/Linfeng-Ye/NCMI.
@@ I Introduction
To this end, following [CMI], we model the DNN-based classification task as a three-state Markov chain, as illustrated in Figure˜1. We quantify the concentration of the output distribution using the conditional mutual information (CMI) between the input and output, conditioned on the ground-truth label, and we quantify separation with $\Gamma$. We then define the normalized conditional mutual information (NCMI) as the ratio between the CMI and $\Gamma$. During training, instead of maximizing the log-likelihood (i.e., minimizing CE), we train the model to minimize its NCMI. After training, we evaluate the model using centroid-based decisions, comparing its outputs to class centroids and via linear probing. Empirically, NCMI outperforms prior state-of-the-art losses and serves as an efficient surrogate for classification. The key contributions are as follows:
@@ II Related work
Within the existing literature, cross-entropy (CE) and its variants are the de facto objectives for classification. Several works have attempted to improve CE. Empirical studies report that DNNs with compact feature clusters usually outperform those with sparse clusters [oyallon2017building, JMLR:v21:20-933, papyan2020prevalence]. These insights have been further analyzed under the Gaussian-mixture assumption on the feature distribution [zarka2021separation]. Building on this line of work, subsequent works augment CE with regularizers to strengthen intra-class concentration and inter-class margins. Specifically, Hui et al. [hui2023cut] add an $\ell_{2}$ penalty to the non-ground-truth entries of the predicted probability distribution, and OPL [OPL] explicitly clusters same-class features while enforcing orthogonality between different classes in the penultimate layer.
Another line of work improves classification accuracy by modifying CE. Focal Loss [lin2017focal] down-weights well-classified examples via a power transformation so training emphasizes hard instances. PolyLoss [leng2022polyloss] reframes standard classification losses as polynomial expansions. Hui et al. propose SquareLoss [hui2021evaluation], and empirically found that squared loss performs on par with or even outperforms CE on modern DNNs. Supervised contrastive learning (SuperCon) [khosla2020supervised] pulls together same-class embeddings and pushes apart different-class embeddings, followed by a linear classifier trained on the frozen features. Further, to mitigate overconfident predictions, label smoothing (LS) [szegedy2016rethinking] softens the one-hot targets, which can inadvertently produce compact class clusters [muller2019does]. AntiClass [kat2025anticlasses] replaces the one-hot target with a one-cold target to mitigate neural collapse.
In contrast, this paper studies the surrogate loss for classification through the lens of information theory. We view a DNN as a mapping from $\boldsymbol{x}\in\mathbb{R}^{d}$ to $\boldsymbol{p}\in\mathcal{P}^{n}$, and we train the model by directly minimizing its conditional mutual information.
@@ III-A Notation
where $\ln$ denotes the logarithm with base $e$, write the CE of the one-hot probability distribution corresponding to $y$ and $\boldsymbol{q}$ as $H(y,\boldsymbol{q})=-\ln\boldsymbol{q}_{y}$. Let $\mathcal{S}$ be a collection of simplex-valued functions on $\mathbb{R}^{n}$, i.e.
@@
$\sigma^{NSF}(\boldsymbol{z})_{j}$ $=\frac{\phi(\boldsymbol{z}_{j})}{\sum_{i=1}^{n}\phi(\boldsymbol{z}_{i})},\penalty 10000\ \penalty 10000\ \phi(\boldsymbol{z}_{i})=\frac{1}{1+e^{\boldsymbol{z}_{i}}};$ (3)
@@ IV-B Implementation and evaluation protocols
Decision based on comparison with centroids. We further evaluate the NCMI-trained model on unseen samples from the test set by comparing them with the centroid of each class. To this end, the NCMI trained model predicts output based on comparison with centroids, specifically, we calculate the KL divergence between its output distribution $\boldsymbol{p}$ of the model and each centroid $\boldsymbol{q}^{v}$ per class $D(\boldsymbol{p}\|\boldsymbol{q}^{v}),\penalty 10000\ v\in[C]$. Then the prediction is made based on the centroid with the smallest KL-divergence value.
@@ V Experiments
To illustrate the effectiveness of NCMI and compare it with some state-of-the-art alternatives, a series of experiments was conducted. Specifically, we conduct experiments on two widely used natural image datasets, namely CIFAR-100 [krizhevsky2009learning] and ImageNet [5206848], as well as two whole-slide image datasets, namely CAMELYON-17 [8447230] and BRACS [brancati2022bracs]. In the tables, NCMI-LP and NCMI-CC denote NCMI evaluated with linear probing (LP) and with centroid comparison (CC), respectively.
@@ V-C Experiments on Whole Slide Image Dataset
Figure 4: The evolution curves of ResNet-18 on CIFAR-100 under all combinations of NSF and feature centering (enabled/disabled). Shown are the epoch-wise trajectories of (a) CMI, (b) $\Gamma$, (c) NCMI, and (d) accuracy.
To assess NCMI beyond natural images, we evaluate on two whole-slide image (WSI) classification benchmarks: CAMELYON-17 [8447230] and BRACS [brancati2022bracs]. CAMELYON-17 comprises 1000 WSIs from five medical centers, providing a diverse and clinically representative cohort. Of these, 500 slides are publicly available with slide-level labels, while the remaining 500 are held out for challenge evaluations. The multi-institutional composition introduces substantial variation in staining and scanning, making CAMELYON-17 a strong test bed for generalization. BRACS is a large-scale WSI dataset curated for breast cancer subtype classification, comprising 547 WSIs collected from several institutions and annotated by expert pathologists into clinically relevant categories: benign tumors, atypical tumors, and malignant tumors. We follow the official split of each dataset into training, validation, and test sets.
@@ VI Ablation Study
To understand how these components affect learning, we visualize the ResNet-18 training trajectories of CMI, $\Gamma$, NCMI, and accuracy on the CIFAR-100 test set in Figure˜4. As shown, the NSF plays a pivotal role in NCMI training: when the softmax function is used to map features to probability distributions, the model fails to converge. Feature centering further stabilizes training. Specifically, with the NSF enabled but without centering, the model collapses around epoch 150; after centering is enabled, it converges properly. Due to the space limit, we further evaluate the effectiveness of NSF and the centering operation by visualizing class clusters and their centers in the appendix.

### v2

@@ Abstract
In this paper, we propose a novel information-theoretic surrogate loss—normalized conditional mutual information (NCMI)—as a drop-in alternative to the de facto cross-entropy (CE) for training deep neural network (DNN)-based classifiers. We first observe that the model’s NCMI is inversely proportional to its accuracy. Building on this insight, we advocate to use NCMI as the surrogate loss for DNN classifier, and propose an alternating algorithm to efficiently minimize the NCMI. Across natural image recognition and whole-slide imaging (WSI) subtyping benchmarks, NCMI-trained models surpass state-of-the-art losses by substantial margins at a computational cost comparable to that of CE. Notably, on ImageNet, NCMI yields a 2.77% top-1 accuracy improvement with ResNet-50 comparing to the CE; on CAMELYON-17, replacing CE with NCMI improves the macro-F1 by 8.6% over the strongest baseline. Gains are consistent across various architectures and batch sizes, suggesting that NCMI is a practical and competitive alternative to CE. All code and data are publicly available at https://github.com/Linfeng-Ye/NCMI.
@@ I Introduction
To this end, following [CMI], we model the DNN-based classification task as a three-state Markov chain, as illustrated in Figure˜1. We quantify the concentration of the output distribution using the conditional mutual information (CMI) between the input and output, conditioned on the ground-truth label, and we quantify separation with $\Gamma$ (see Section˜III-B). We then define the normalized conditional mutual information (NCMI) as the ratio between the CMI and $\Gamma$. During training, instead of maximizing the log-likelihood (i.e., minimizing CE), we train the model to minimize its NCMI. After training, we evaluate the model using centroid-based decisions by comparing its outputs to class centroids, and via linear probing. Empirically, NCMI outperforms prior state-of-the-art losses and serves as an efficient surrogate for classification. The key contributions are as follows:
@@ II Related work
Within the existing literature, CE and its variants are the de facto objectives for classification. Several works have attempted to improve CE. Empirical studies report that DNNs with compact feature clusters usually outperform those with sparse clusters [oyallon2017building, JMLR:v21:20-933, papyan2020prevalence]. These insights have been further analyzed under the Gaussian-mixture assumption on the feature distribution [zarka2021separation]. Building on this line of work, subsequent works augment CE with regularizers. Specifically, Hui et al. [hui2023cut] add an $\ell_{2}$ penalty to the non-ground-truth entries of the predicted probability distribution, and OPL [OPL] explicitly clusters same-class features while enforcing orthogonality between different classes in the penultimate layer.
Another line of work improves classification accuracy by modifying CE. Focal Loss [lin2017focal] down-weights well-classified examples via a power transformation so training emphasizes hard instances. PolyLoss [leng2022polyloss] reframes standard classification losses as polynomial expansions. Hui et al. propose SquareLoss [hui2021evaluation], and empirically found that squared loss performs on par with or even outperforms CE on modern DNNs. Supervised contrastive learning (SupCon) [khosla2020supervised] pulls together same-class embeddings and pushes apart different-class embeddings, followed by a linear classifier trained on the frozen features. Further, to mitigate overconfident predictions, label smoothing (LS) [szegedy2016rethinking] softens the one-hot targets, which can inadvertently produce compact class clusters [muller2019does]. AntiClass [kat2025anticlasses] replaces the one-hot target with a one-cold target to mitigate neural collapse.
In contrast, this paper studies the surrogate loss for classification through the lens of information geometry. We view a DNN as a mapping from $\boldsymbol{x}\in\mathbb{R}^{d}$ to $\boldsymbol{p}\in\mathcal{P}^{n}$. NCMI trains the DNN by encouraging the intra-class concentration and inter-class separation of the output distribution cluster.
@@ III-A Notation
where $\ln$ denotes the logarithm with base $e$, write the CE of the one-hot probability distribution corresponding to $y$ and $\boldsymbol{q}$ as $H(y,\boldsymbol{q})=-\ln\boldsymbol{q}_{y}$. Let $\mathcal{S}$ be a collection of probability-simplex-valued functions on $\mathbb{R}^{n}$, i.e.
@@
$\sigma^{NSF}(\boldsymbol{z})_{j}$ $=\frac{\phi(\boldsymbol{z}_{j})}{\sum_{i=1}^{n}\phi(\boldsymbol{z}_{i})},\penalty 10000\ \penalty 10000\ \phi(\boldsymbol{z}_{i})=\frac{1}{1+e^{-\boldsymbol{z}_{i}}};$ (3)
@@ IV-B Implementation and evaluation protocols
Decision based on centroids comparison. We further evaluate the NCMI-trained model on unseen samples from the test set by comparing them with the centroid of each class. To this end, the NCMI trained model predicts output based on comparison with centroids, specifically, we calculate the KL divergence between its output distribution $\boldsymbol{p}$ of the model and each centroid $\boldsymbol{q}^{v}$ per class $D(\boldsymbol{p}\|\boldsymbol{q}^{v}),\penalty 10000\ v\in[C]$. Then the prediction is made based on the centroid with the smallest KL-divergence value.
@@ V Experiments
To illustrate the effectiveness of NCMI and compare it with some state-of-the-art alternatives, a series of experiments was conducted. Specifically, we conduct experiments on two widely used natural image datasets, namely CIFAR-100 [krizhevsky2009learning] and ImageNet [5206848], as well as two whole-slide image datasets, namely CAMELYON-17 [8447230] and BRACS [brancati2022bracs]. In the tables, NCMI-LP and NCMI-CC denote NCMI evaluated with linear probing (LP) and with centroids comparison (CC), respectively.
@@ V-C Experiments on Whole Slide Image Dataset
Figure 4: The evolution curves of ResNet-18 on CIFAR-100 under all combinations of NSF and feature centering (enabled/disabled), w.o.C/w.C denote without/with centering, and SM/NSF denote applying softmax/NSF. Shown are the epoch-wise trajectories of (a) CMI, (b) $\Gamma$, (c) NCMI, and (d) accuracy.
To assess NCMI beyond natural image datasets, we evaluate on two whole-slide image (WSI) classification benchmarks: CAMELYON-17 [8447230] and BRACS [brancati2022bracs]. CAMELYON-17 comprises 1000 WSIs from five medical centers, providing a diverse and clinically representative cohort. Of these, 500 slides are publicly available with slide-level labels, while the remaining 500 are held out for challenge evaluations. The multi-institutional composition introduces substantial variation in staining and scanning, making CAMELYON-17 a strong test bed for generalization. BRACS is a large-scale WSI dataset curated for breast cancer subtype classification, comprising 547 WSIs collected from several institutions and annotated by expert pathologists into clinically relevant categories: benign tumors, atypical tumors, and malignant tumors. We follow the official split of each dataset into training, validation, and test sets.
@@ VI Ablation Study
To understand how these components affect learning, we visualize the ResNet-18 training trajectories of CMI, $\Gamma$, NCMI, and accuracy on the CIFAR-100 test set in Figure˜4. As shown, the NSF plays a pivotal role in NCMI training: when the softmax function is used to map features to probability distributions, the model fails to converge. Feature centering further stabilizes training. Specifically, with the NSF enabled but without centering, the model collapses around epoch 150; after centering is enabled, it converges properly. Due to the space limit, we further evaluate the effectiveness of NSF and the centering operation by visualizing class clusters and their centers in the appendix [^2] Extended version (with appendix): https://arxiv.org/pdf/2601.02543..